# MT3DNet: Multi-Task learning Network for 3D Surgical Scene Reconstruction

### v1

@@ 1 Introduction
The progress in comprehending surgical scenes heavily relies on the EndoVis2018 datasets, which feature annotated robotic instrument data for tasks like segmentation and detection. Despite containing seven tube-like instruments, these datasets lack monocular depth annotations for MIS settings. To address this, the authors employed the state-of-the-art STL depth estimation model, Depth Anything, to generate depth maps for training the MTL model.
@@ 3 Methodology
3.1 MTL Architecture
The proposed model architecture comprises an Encoder 3.2, Decoder 3.2.1, and task-specific heads 3.2.2, integrated into a unified Multi-Task Learning Transformer (MT3DNet) 3.2.3 framework is depicted in Fig. 1.
@@ 4.2 Adversarial Weight Updates for Multi-Task Learning
This adversarial adjustment is then integrated with gradient-based updates. Task gradients are collected and their interrelations are captured using a gradient Gram matrix. We solve for optimal weights by employing a regularized MWU-inspired approach:
@@
The use of these weight updates drives this multi-task learning framework to reliably align gradients and optimise the tasks without explicit optimisation dynamics, hence being robust and efficient in the overall learning process.
@@ 5 Experimental Results
In our experiment, we used the Adan optimizer [23] with an initial learning rate of $0.001$ and a weighted decay of $0.02$. We also incorporated a Reduce on Plateau scheduler and an exponential learning rate scheduler to enhance learning dynamics, ensuring efficient convergence and improved performance throughout training, contributing to robustness and effectiveness. Leveraging the publicly available EndoVis18 dataset, comprising 14 videos of porcine surgeries with The Da Vinci Xi [10] surgical system, provided meticulous annotations for the entire surgical scene and individual instrument parts. This comprehensive dataset provides the necessary ground truth for segmentation and detection annotations. However, it was missing depth annotations, which are essential for learning depth data.
@@ 6 Conclusion
In this study, a novel multi-task learning framework called MT3DNet was proposed for segmenting, localizing, and estimating depth in image-assisted Minimally Invasive Surgery environments. With the use of Adversarial Weight Update, challenges of optimization inherent with the management of multiple tasks simultaneously were addressed to bring about optimality. Extensive experiments on the EndoVis2018 dataset demonstrated the model’s ability to handle all three tasks, which demonstrated the effectiveness of our methodologies Importantly, the approach enables 3D reconstruction by integrating segmentation, depth estimation, and object detection, which improves surgical scene understanding significantly over prior studies that lack 3D capabilities. The novel approach with the prototype shows potential for use in diverse scenarios with segmentation, detection, and depth annotation. Further research should be done in advancing 3D reconstruction across various different avenues using this approach.

### v2

@@ Abstract
\$*\$\$*\$footnotetext: These authors contributed equally to this work
@@ 1 Introduction
The progress in comprehending surgical scenes heavily relies on the EndoVis2018 datasets, which feature annotated robotic instrument data for tasks like segmentation and detection. Despite containing seven tube-like instruments, these datasets lack monocular depth annotations for MIS settings. To address this, the authors employed the state-of-the-art STL depth estimation model, Depth Anything[24], to generate depth maps for training the MTL model.
@@ 3 Methodology
The proposed model architecture comprises an Encoder, Decoder, and task-specific heads, integrated into a unified Multi-Task Learning Transformer (MT3DNet) framework is depicted in Fig. 1.
@@ 4.2 Adversarial Weight Updates for Multi-Task Learning
To further illustrate this, the adversarial updates of the weights make use of loss values across tasks, driving optimization. It adaptively amplifies the higher loss tasks so that low performers get more attention in the training process. Details for this are available in Algorithm 1.
@@
The optimal weight computation method accompanies the adversarial update method. This method lines up task gradients to diminish inter-task conflicts and stabilize learning dynamics. The gradient alignment method makes sure that the learning dynamics are robust against any kind of change in the task complexity. Algorithm 2 describes the procedure.
@@
The use of these weight updates drives this multi-task learning framework to reliably align gradients and optimize the tasks without explicit optimization dynamics, hence being robust and efficient in the overall learning process.
@@ 5 Experimental Results
In our experiment, we used the Adan optimizer [23] with an initial learning rate of $0.001$ and a weighted decay of $0.02$. We also incorporated a Reduce on Plateau scheduler and an exponential learning rate scheduler to enhance learning dynamics, ensuring efficient convergence and improved performance throughout training, contributing to robustness and effectiveness. Leveraging the publicly available EndoVis18 dataset, comprising 14 videos of porcine surgeries with The Da Vinci Xi [10] surgical system, provided meticulous annotations for the entire surgical scene and individual instrument parts. This comprehensive dataset provides the necessary ground truth for segmentation and detection annotations. However, it was missing depth annotations, which are essential for learning depth data. To address this, depth maps were created using the state-of-the-art STL depth estimation model, Depth Anything[24], which takes in visual input to infer pixel-wise depth, thus creating high-quality synthetic depth annotations optimized for multi-task learning.
@@ 6 Conclusion
In this study, a novel multi-task learning framework called MT3DNet was proposed for segmenting, localizing, and estimating depth in image-assisted Minimally Invasive Surgery environments. With the use of Adversarial Weight Update, challenges of optimization inherent with the management of multiple tasks simultaneously were addressed to bring about optimality. Extensive experiments on the EndoVis2018 dataset demonstrated the model’s ability to handle all three tasks, which demonstrated the effectiveness of our methodologies. Importantly, the approach enables 3D reconstruction by integrating segmentation, depth estimation, and object detection, which improves surgical scene understanding significantly over prior studies that lack 3D capabilities. The novel approach with the prototype shows potential for use in diverse scenarios with segmentation, detection, and depth annotation. Further research should be done in advancing 3D reconstruction across various different avenues using this approach.